OpenStudy (anonymous):

Write the equation using function notation. 4x+3y=12 Can someone show me how to do this?

hero (hero):

Solve for y, then replace y with f(x)

OpenStudy (anonymous):

which is independent var, which is dependent var

hero (hero):

y = (12 - 4x)/3 f(x) = (12 - 4x)/3

OpenStudy (anonymous):

divide the 12 and -4x by 3?